OpenStudy (anonymous):

What is the solution to the system of equations that contains −3x + y = 3 and 2x − y = −1? a.no solution b.(−1, 3) c.infinite number of solutions d.(−2, −3)

OpenStudy (anonymous):

(-2, -3)

OpenStudy (anonymous):

you sure ? this is really important

OpenStudy (anonymous):

Positive. Would you like me to explain how I got my answer?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

y=-3x+3 y=2x+1

OpenStudy (anonymous):

Graph it starting at 3 and have the slope of 3 then start at 1 with a slope of 2

OpenStudy (anonymous):

ok , can you help me with three other questions ?

OpenStudy (anonymous):

I will do my best

OpenStudy (anonymous):

The point (2, 1) is a solution to which of the following systems of equations? 2x + y = 5 and −2x + y = 2 −2x − y = −5 and 2x − y = 2 −x + y = 3 and −3x + y = −5 x + y = 3 and 3x − y = 5

OpenStudy (anonymous):

D X+Y+3 and 3X-Y=5

OpenStudy (anonymous):

Anymore questions?

OpenStudy (anonymous):

On a graph, the solution to a system of linear equations is represented by __________. the x–intercept of each line the y–intercept of each line all the points on either line the intersection of the lines

OpenStudy (anonymous):

The intersection of the lines

OpenStudy (anonymous):

last question :) Which of the following values for c would mean that the system of equations 2x − 3y = 1 and cx − 3y = 2 would not have any solutions? 2 3 1 0

OpenStudy (anonymous):

Easy 2

OpenStudy (anonymous):

thanks ! you just saved my life

OpenStudy (anonymous):

:D Let me know what you get!

OpenStudy (anonymous):

100% !!